Albert bought his Bycicle for $275. He sold it for $350.
A.) What was the amount of his profit
B.) What was the amount of his percentage profit.



Please show all working👨‍💻 please🙏🙏

Answer by math_tutor2020(3835)


Part A

profit = revenue - cost
profit = 350 - 275
profit = 75

====================================================
Part B

Profit percentage = (profit/cost)*100%
Profit percentage = (75/275)*100%
Profit percentage = (0.2727)*100%
Profit percentage = 27.27%

The value is approximate.
This says that roughly 27.27% of the cost is profit.